OpenStudy (anonymous):

A newspaper reports the following changes in the temperature of a city over 4 days: 1.34 degrees Celsius, negative 5 over 7 degrees Celsius, −0.75 degrees Celsius, 4 over 9 degrees Celsius Write an expression to show the average daily change in the temperature of the city. Write the steps to solve the expression. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

ill give medal and fan for answer pls help

OpenStudy (fibonaccichick666):

why did you reopen it without responding?

OpenStudy (anonymous):

@DangerousJesse

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

help pls

OpenStudy (fibonaccichick666):

http://openstudy.com/study#/updates/53f5327ce4b0f0e909e00974

OpenStudy (anonymous):

m8 can u help

OpenStudy (aum):

There are 4 temperatures given. To find the average, add the 4 numbers and divide by 4. "Write an expression to show the average daily change in the temperature of the city." The expression is: numerator: temperature 1 + temperature 2 + temperature 3 + temperature 4. Denominator: 4 (just substitute the values for each temperature.)

OpenStudy (anonymous):

oh ok thx

OpenStudy (aum):

For part b) since they want the answer in decimal rounded to two decimal places, convert each fraction in the numerator to decimal, add the numbers and divide by 4. Round the answer to two decimal places.

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} x&y \\\hline\\ {\color{red}{ 1}}&{\color{blue}{ 1.34}}\\ 2&-\frac{5}{7}\\ 3&-0.75\\ {\color{red}{ 4}}&{\color{blue}{ \frac{4}{9}}} \end{array}\qquad slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\\\ \quad \\\\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} }\)

OpenStudy (jdoe0001):

the slope will give you the "average rate of change" thus

OpenStudy (aum):

a) Average daily change in the temperature = \[\frac{1.34 + (-5/7) + (-0.75) + 4/9}{4} = \frac{1.34 - 5/7 - 0.75 + 4/9}{4} \]b) Convert fractions to decimals: Average daily change in the temperature = \[\frac{1.34 + (-5/7) + (-0.75) + 4/9}{4} = \frac{1.34 - 0.71 - 0.75 + 0.44}{4} = \frac{0.32}{4} = 0.08 \]